# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: origins-of-hinduism Description: Many historians argue that Hinduism started to develop when Indus Valley and Aryan cultures mixed. Look at the following information about some common Hindu beliefs and religious practices, or actions, that developed over time. Then use the information to answer the question below. Some Hindus believe that there are hundreds of different gods, but some think there is just one god who has many different characteristics. Hindus in different areas often celebrate different holidays and festivals. Based on the information above, which of the following statements is true? Choices: [(a) There are many different ways of practicing Hinduism., (b) There is only one main way of practicing Hinduism.] Answer index: 0
Subject: Social Studies Skill: origins-of-buddhism Description: Historians agree that Buddhism began with the teachings of Siddhartha Gautama (si-DAR-tuh GOW-tuh-muh). During his life, Siddhartha became a great teacher and was given the title "buddha" (BOO-duh). Read about some common Buddhist beliefs. Then answer the question below. There were buddhas who lived before Siddhartha, and there will be buddhas who come after him. Siddhartha is often called "the Buddha," instead of just "buddha," because he is the buddha of this age. What does it mean to say that Siddhartha is "the buddha of this age"? Choices: [(a) He was the buddha who lived most recently., (b) He was not very old when he became a buddha., (c) He was less important than future buddhas., (d) He was the first buddha who ever lived.] Answer index: 0
Subject: Social Studies Skill: the-ghana-empire Description: Historians are not sure who founded the ancient Kingdom of Ghana. However, legends about Ghana’s first king have been passed down by griots, or people who tell stories about West African history. Read a story that is told by griots. Then complete the text below. There were two brothers named Dyabe and Khine, and their father was the leader of their tribe. At first, their father wanted Khine to be the next leader. But Dyabe showed he was more clever than Khine, and their father changed his mind. Khine grew jealous and tried to have Dyabe killed, so Dyabe fled to the wilderness to save himself. Out in the wilderness, Dyabe met the leaders of four different armies. He convinced these army leaders to make him their ruler. Once he had the support of four armies, Dyabe decided to build a kingdom. But first, he needed to find a capital city for his kingdom. Dyabe traveled to a city that was guarded by the snake-god Bida. Bida agreed to let Dyabe use the city as his capital in return for a yearly sacrifice. So, Dyabe built the capital and created the Kingdom of Ghana. Complete the text. According to this story, the Kingdom of Ghana was founded by a man named ___. He created the kingdom after making a deal with ___. Choices: [(a) Dyabe . . . Khine, (b) Khine . . . the snake-god Bida, (c) Dyabe . . . the snake-god Bida, (d) Khine . . . Dyabe] Answer index: 2
@@
Subject: Social Studies Skill: the-middle-ages-after-1100 Description: As kingdoms in Western Europe became larger, kings gained more wealth and power. Some kings began to take advantage of this power and abused the people in their kingdoms. In some parts of Europe, people responded to this behavior by trying to limit the power of kings. Read the passage about one king who abused his power. Then complete the sentence below. King John ruled England from 1199 to 1216 CE. In 1215, a group of lords demanded that King John end an unfair tax that was against the law. King John refused to end the tax, so the lords captured the city of London, where King John ruled. After capturing the city, the lords forced King John to sign an agreement that would limit his power as the king. This agreement became known as the Magna Carta, or Great Charter. The Magna Carta was an agreement that ___ in England. Choices: [(a) began a long and expensive war, (b) overthrew and killed the king, (c) increased taxes on the lords, (d) controlled the power of King John] Answer index: 3
Subject: Social Studies Skill: feudal-europe Description: By the early Middle Ages, many different Western European landowners had started organizing society in a new way. Historians call this new way of organizing society feudalism. Read the passage about how feudalism began. Then complete the text. After the fall of the Roman Empire, the Roman government no longer controlled Western Europe. So, individuals fought each other for control of land and the people living on it. The winners gained wealth and power, and people living on the land were forced to work for them. Some new landowners became powerful and started calling themselves kings, but they couldn’t manage all their land on their own. So, they divided their land up and gave small sections to loyal followers. These loyal followers were called lords, and they helped to control the land and people for the king. Using the information in the passage, complete the text. Feudalism is a system where ___. Choices: [(a) people voted to elect leaders and decide who would own the land, (b) land and people were controlled by lords who were loyal to a king, (c) every member of society was given an equal amount of land to farm by their leader] Answer index: 1
@@
Subject: Social Studies Skill: classical-athens-government-and-culture Description: Not all decisions in classical Athens were made by all the citizens. Instead, a full-time council of citizens made many of the everyday political decisions in the city. Here is a list of some facts about the Athenian council: This Athenian council was made up of 500 citizens. A councilman could only serve for one year at a time. Which of the following statements might come from an Athenian complaining about these rules? Choices: [(a) I think that because the selection is random, any foolish person can become a councilman., (b) Last time I was on the council, I was impressed with how many people had good ideas., (c) I’m glad that 500 people share power, since it means no single person can take over the city.] Answer index: 0
@@
Subject: Social Studies Skill: caesar-and-the-end-of-the-roman-republic Description: Pompey agreed to help the Senate and try to prevent Caesar from becoming more powerful. Read the passage. Then answer the question below. In January of 49 BCE, Caesar was with his army far away from Rome. Pompey wrote to Caesar and demanded that Caesar give up his army and return to Rome. Caesar faced an important choice. On the one hand, he could obey Pompey and give up his army. But he knew the Senate didn’t like him, and he worried that he would be arrested. On the other hand, he could refuse to give up his army. But then the Senate would probably think that Caesar wanted to start a war. After thinking about it for a long time, Caesar chose the second option. He decided to attack Pompey and the Senate before they could raise an army to stop him. Based on the passage above, how did Caesar respond to Pompey’s demands? Choices: [(a) He disobeyed Pompey but offered to cooperate with the Senate., (b) He disobeyed Pompey and started a war against Pompey and the Senate., (c) He decided to give up his army and return to Rome to face punishment.] Answer index: 1
Subject: Social Studies Skill: roman-government-monarchy-to-republic Description: According to present-day historians, the Roman Republic lasted from 509 BCE to about 27 BCE. The word republic comes from the ancient Roman term res publica. Look at the meaning of the term res publica. Then answer the question below. res publica: a community committed to following the same rules and protecting each other’s things and families for the good of all Based on the meaning above, which of the following statements describe something similar to what Romans meant by res publica? Choices: [(a) a group of friends that do whatever they want, whenever they want, (b) a class project group where everyone follows the same guidelines and benefits equally from working together, (c) a soccer team where bad players have to do what the coach says, but the best players don’t] Answer index: 1
Subject: Social Studies Skill: french-and-english-expeditions-part-2 Description: After John Cabot’s expedition, the English government did not sponsor another expedition to the Americas for more than 50 years. Then, in the mid-1500s, the English government began sending expeditions again. Read the passage about the events in England that created new interest in expeditions. Then answer the question below. In the middle of the 1500s, England’s main source of wealth was trading English cloth in foreign markets. Thousands of people in England made their livings by producing or selling cloth, and the English government made money by taxing the sale of cloth. But in 1551, some foreign powers stopped allowing people to purchase English cloth. Many English people who worked in the cloth industry lost their jobs, and the government could not raise as much money. These conditions created a national crisis. The Queen of England, Elizabeth I, needed to find new sources of wealth to address the crisis. Queen Elizabeth thought that sending expeditions to the Americas might be one way to help. Based on the passage, why did Queen Elizabeth I decide to begin sending expeditions to the Americas? Choices: [(a) She wanted to stop Spain from purchasing English cloth., (b) She needed to find a new way to make money for England after the cloth crisis., (c) She was curious to learn more about what the Americas looked like.] Answer index: 1
@@
Subject: Social Studies Skill: comparing-athens-and-sparta-part Description: Although city-states such as Athens could trade for food, Sparta was surrounded by mountains that made trade difficult. So, the Spartans relied on agriculture. The ancient Greek historian Pausanias described the role of helots, or enslaved people, in Spartan agriculture. Read the passage. Then complete the text below. The helots had to bring half of all the crops they grew to the Spartans. Helots who disobeyed were severely punished. The poet Tyrtaeus compared the helots to donkeys exhausted by a heavy load, under painful duty to bring their Spartan masters half of the crops from the land they ploughed and cared for. Adapted from Pausanias, Description of Greece, translated by W.H.S. Jones and H.A. Ormerod. Complete the text. The helots were enslaved people in Sparta. The Spartans relied on the helots ___. Choices: [(a) to educate their children, (b) to farm their land and raise crops, (c) to trade what was farmed, (d) to train their children to farm] Answer index: 1
@@
Subject: Social Studies Skill: thomas-edison Description: Several inventors worked on lightbulbs before Thomas Edison did. How were Thomas Edison’s lightbulbs different? Choices: [(a) His lightbulbs were free., (b) His lightbulbs were made from all recycled materials., (c) His lightbulbs could last for several weeks or months., (d) His lightbulbs came in different colors.] Answer index: 2
@@
Subject: Social Studies Skill: the-revolutionary-war-turning-the-tide-of-war Description: Both the Continental Army and people on the homefront had trouble paying for the things that they needed during the war. Why? Choices: [(a) Inflation caused prices to rise., (b) The Continental Congress stopped printing Continentals., (c) The British took gold reserves from the Thirteen Colonies., (d) Most of the money was spent on military supplies.] Answer index: 0
@@
Subject: Social Studies Skill: foundations-of-maya-civilization Description: Complete the text. Archaeologists believe that the ancestors of the Maya first arrived in Central America more than 12,000 years ago. At first, these people were ___ and traveled long distances to find food. But over time, they learned to grow crops instead. By at least 1500 BCE, they began to settle in permanent ___. This was the beginning of the Maya civilization. Choices: [(a) explorers . . . cities, (b) farmers . . . palaces, (c) hunter-gatherers . . . villages] Answer index: 2
Subject: Social Studies Skill: classical-athens-geography-and-society Description: Complete the text. For most of its history, ancient Athens ruled itself and was not part of a larger country. In other words, Athens was a ___. Choices: [(a) region, (b) city-state, (c) kingdom] Answer index: 1
Subject: Social Studies Skill: the-mongol-empire Description: The founder of the Mongol Empire was a man named Temujin. He was born in Mongolia in the 1100s. The passage below comes from a Mongolian history text. In the passage, an older person is telling Temujin what Mongolia was like before he was born. Read the passage. Then answer the question below. Before you were born, the stars turned in the heavens. Everyone was fighting. Instead of sleeping, they robbed each other of their possessions. . . . Instead of resting, they fought each other. In such a world one did not live as one wished, but rather in constant conflict. There was no rest, only battle. There was no affection, only mutual slaughter. Adapted from The Secret History of the Mongols, paragraph 254 Based on the passage, which choice best summarizes what Mongolian society was like in the 1100s? Choices: [(a) Mongolian society was peaceful, but foreign invaders often attacked Mongolia. As a result, Mongols were often forced to fight, and many Mongols died., (b) People in Mongolia often fought wars with China, their neighbor. As a result, people often went to China to fight., (c) People in Mongolia were fighting with one another and stealing each others’ goods. As a result, people could not live peaceful lives.] Answer index: 2
Subject: Social Studies Skill: foundations-of-aztec-civilization Description: Today, the part of Central America where the Aztec Empire existed is known as Mesoamerica. There are two ways of defining a region like Mesoamerica: using a geographic definition, which focuses on an area of land and its features, such as mountain ranges and bodies of water using a cultural definition, which focuses on the customs and practices that people living there share Which of the following statements uses a geographic definition of Mesoamerica? Choices: [(a) Mesoamerica is a place where people used a 260-day calendar., (b) Mesoamerica is a place in Central America where people built religious temples in the form of pyramids., (c) Mesoamerica is a region of Central America that includes the Valley of Mexico and the Yucatan Peninsula.] Answer index: 2
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z Description: Pick the uppercase letter that matches. k Choices: [(a) T, (b) R, (c) K] Answer index: 2
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z Description: Pick the lowercase letter that matches. C Choices: [(a) c, (b) v, (c) o] Answer index: 0
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r Description: Pick the uppercase letter that matches. a Choices: [(a) N, (b) A] Answer index: 1
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r Description: Pick the lowercase letter that matches. A Choices: [(a) u, (b) a, (c) n] Answer index: 1
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y Description: Pick the lowercase letter that matches. L Choices: [(a) l, (b) k, (c) m] Answer index: 0
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y Description: Pick the uppercase letter that matches. m Choices: [(a) H, (b) J, (c) M] Answer index: 2
@@
Subject: Language Arts Skill: distinguish-characters-points-of-view Description: The following texts both describe a school bus ride on the first day of school. This text is told from the point of view of Jeffrey, a student going to a new school: I felt ready for my first shift as a school bus driver. I turned the long bus at each corner with ease. My cheerful mood made the students smile. I was proud of myself as the bus approached the brick school building. My first trip had gone well! How is Ms. Romero’s point of view different from Jeffrey’s? Choices: [(a) Ms. Romero describes feeling uneasy about a new experience., (b) Ms. Romero expresses a feeling of pride about the first day of school.] Answer index: 1
@@
Subject: Language Arts Skill: determine-the-meanings-of-words-with-greek-and-latin-roots Description: What does the word autopilot mean? Choices: [(a) a tool that pilots use when they want to make the airplane fly faster, (b) a device that flies an airplane on its own, without a pilot, (c) a map or drawing that shows an airplane’s route] Answer index: 1
Subject: Language Arts Skill: describe-the-difference-between-related-words Description: What is the difference between being vicious and being mean? Choices: [(a) being vicious is worse, (b) being mean shows more anger] Answer index: 0
Subject: Language Arts Skill: which-word-does-not-rhyme Description: Which word does not rhyme? Choices: [(a) pack, (b) pick, (c) sick] Answer index: 0
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word Description: Complete the sentence. ___ pig is big. Choices: [(a) That, (b) Went] Answer index: 0
Subject: Language Arts Skill: to-have-use-the-correct-form Description: Complete the sentence. Mrs. Barton’s rug ___ blue flowers on it. Choices: [(a) has, (b) have] Answer index: 0
Subject: Language Arts Skill: complete-the-sentence-with-a-two-syllable-word Description: Which word makes more sense in the sentence? We ride in the ___ Choices: [(a) wagon, (b) robin] Answer index: 0
Subject: Language Arts Skill: complete-the-sentence-with-a-three-letter-consonant-blend-word Description: Which word makes more sense in the sentence? Look, that big fish is swimming up the ___ Choices: [(a) stream, (b) strict] Answer index: 0
Subject: Language Arts Skill: use-subordinating-conjunctions Description: Complete the sentence with the better conjunction. ___ Joseph arrives soon, we won’t get to the movie on time. Choices: [(a) Unless, (b) After] Answer index: 0
@@
Subject: Language Arts Skill: use-the-correct-homophone Description: Complete the sentence with the correct homophone. Given that there were more than one hundred guests at the wedding, how many ___ do you think Martina and Cooper have received? Choices: [(a) presence, (b) presents] Answer index: 1
Subject: Language Arts Skill: use-words-as-clues-to-the-meanings-of-greek-and-latin-roots Description: The word millimeter contains the root mille. What does the root mille mean? Choices: [(a) measure, (b) one hundred, (c) one thousand] Answer index: 2
Subject: Language Arts Skill: analogies Description: Complete the analogy. Ohio is to United States as Earth is to Choices: [(a) moon, (b) solar system] Answer index: 1
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-or-a-fragment Description: Is this a complete sentence? Monica speaks four languages. Choices: [(a) Monica speaks four languages.] Answer index: 0
Subject: Language Arts Skill: is-it-a-phrase-or-a-clause Description: Is the group of words in bold a phrase or a clause? **Emily’s suitcase was too large to fit in the overhead bin**, so unfortunately she had to check it. Choices: [(a) phrase, (b) clause] Answer index: 1
@@
Subject: Language Arts Skill: choose-between-subject-and-object-pronouns Description: Select the correct pronoun to complete the sentence. "Thanks to your efforts, ___ have emerged as leaders in the energy industry," Mr. Bullock announced at the company meeting. Choices: [(a) us, (b) we] Answer index: 1

### v3

@@ Appendix D Summary of Skills
Description: Many historians argue that Hinduism started to develop when Indus Valley and Aryan cultures mixed. Look at the following information about some common Hindu beliefs and religious practices, or actions, that developed over time. Then use the information to answer the question below.
Some Hindus believe that there are hundreds of different gods, but some think there is just one god who has many different characteristics.
Hindus in different areas often celebrate different holidays and festivals.
Based on the information above, which of the following statements is true?
Choices: [(a) There are many different ways of practicing Hinduism., (b) There is only one main way of practicing Hinduism.]
Answer index: 0
Subject: Social Studies
Skill: origins-of-buddhism
Description: Historians agree that Buddhism began with the teachings of Siddhartha Gautama (si-DAR-tuh GOW-tuh-muh). During his life, Siddhartha became a great teacher and was given the title "buddha" (BOO-duh).
Read about some common Buddhist beliefs. Then answer the question below.
There were buddhas who lived before Siddhartha, and there will be buddhas who come after him.
Siddhartha is often called "the Buddha," instead of just "buddha," because he is the buddha of this age.
What does it mean to say that Siddhartha is "the buddha of this age"?
Choices: [(a) He was the buddha who lived most recently., (b) He was not very old when he became a buddha., (c) He was less important than future buddhas., (d) He was the first buddha who ever lived.]
Answer index: 0
Subject: Social Studies
Skill: the-ghana-empire
Description: Historians are not sure who founded the ancient Kingdom of Ghana. However, legends about Ghana’s first king have been passed down by griots, or people who tell stories about West African history.
Read a story that is told by griots. Then complete the text below.
There were two brothers named Dyabe and Khine, and their father was the leader of their tribe. At first, their father wanted Khine to be the next leader. But Dyabe showed he was more clever than Khine, and their father changed his mind. Khine grew jealous and tried to have Dyabe killed, so Dyabe fled to the wilderness to save himself.
Out in the wilderness, Dyabe met the leaders of four different armies. He convinced these army leaders to make him their ruler. Once he had the support of four armies, Dyabe decided to build a kingdom. But first, he needed to find a capital city for his kingdom. Dyabe traveled to a city that was guarded by the snake-god Bida. Bida agreed to let Dyabe use the city as his capital in return for a yearly sacrifice. So, Dyabe built the capital and created the Kingdom of Ghana.
Complete the text.
According to this story, the Kingdom of Ghana was founded by a man named ___. He created the kingdom after making a deal with ___.
Choices: [(a) Dyabe . . . Khine, (b) Khine . . . the snake-god Bida, (c) Dyabe . . . the snake-god Bida, (d) Khine . . . Dyabe]
Answer index: 2
Subject: Social Studies
Skill: the-middle-ages-after-1100
@@
Description: As kingdoms in Western Europe became larger, kings gained more wealth and power. Some kings began to take advantage of this power and abused the people in their kingdoms. In some parts of Europe, people responded to this behavior by trying to limit the power of kings. Read the passage about one king who abused his power. Then complete the sentence below.
King John ruled England from 1199 to 1216 CE. In 1215, a group of lords demanded that King John end an unfair tax that was against the law. King John refused to end the tax, so the lords captured the city of London, where King John ruled. After capturing the city, the lords forced King John to sign an agreement that would limit his power as the king. This agreement became known as the Magna Carta, or Great Charter.
The Magna Carta was an agreement that ___ in England.
Choices: [(a) began a long and expensive war, (b) overthrew and killed the king, (c) increased taxes on the lords, (d) controlled the power of King John]
Answer index: 3
Subject: Social Studies
Skill: feudal-europe
Description: By the early Middle Ages, many different Western European landowners had started organizing society in a new way. Historians call this new way of organizing society feudalism.
Read the passage about how feudalism began. Then complete the text.
After the fall of the Roman Empire, the Roman government no longer controlled Western Europe. So, individuals fought each other for control of land and the people living on it. The winners gained wealth and power, and people living on the land were forced to work for them.
Some new landowners became powerful and started calling themselves kings, but they couldn’t manage all their land on their own. So, they divided their land up and gave small sections to loyal followers. These loyal followers were called lords, and they helped to control the land and people for the king.
Using the information in the passage, complete the text.
Feudalism is a system where ___.
Choices: [(a) people voted to elect leaders and decide who would own the land, (b) land and people were controlled by lords who were loyal to a king, (c) every member of society was given an equal amount of land to farm by their leader]
Answer index: 1
Subject: Social Studies
Skill: classical-athens-government-and-culture
Description: Not all decisions in classical Athens were made by all the citizens. Instead, a full-time council of citizens made many of the everyday political decisions in the city.
Here is a list of some facts about the Athenian council:
This Athenian council was made up of 500 citizens.
A councilman could only serve for one year at a time.
Which of the following statements might come from an Athenian complaining about these rules?
@@
Choices: [(a) I think that because the selection is random, any foolish person can become a councilman., (b) Last time I was on the council, I was impressed with how many people had good ideas., (c) I’m glad that 500 people share power, since it means no single person can take over the city.]
Answer index: 0
Subject: Social Studies
Skill: caesar-and-the-end-of-the-roman-republic
Description: Pompey agreed to help the Senate and try to prevent Caesar from becoming more powerful. Read the passage. Then answer the question below.
In January of 49 BCE, Caesar was with his army far away from Rome. Pompey wrote to Caesar and demanded that Caesar give up his army and return to Rome.
@@
Caesar faced an important choice. On the one hand, he could obey Pompey and give up his army. But he knew the Senate didn’t like him, and he worried that he would be arrested. On the other hand, he could refuse to give up his army. But then the Senate would probably think that Caesar wanted to start a war. After thinking about it for a long time, Caesar chose the second option. He decided to attack Pompey and the Senate before they could raise an army to stop him.
Based on the passage above, how did Caesar respond to Pompey’s demands?
Choices: [(a) He disobeyed Pompey but offered to cooperate with the Senate., (b) He disobeyed Pompey and started a war against Pompey and the Senate., (c) He decided to give up his army and return to Rome to face punishment.]
Answer index: 1
Subject: Social Studies
Skill: roman-government-monarchy-to-republic
Description: According to present-day historians, the Roman Republic lasted from 509 BCE to about 27 BCE. The word republic comes from the ancient Roman term res publica. Look at the meaning of the term res publica. Then answer the question below.
res publica: a community committed to following the same rules and protecting each other’s things and families for the good of all
Based on the meaning above, which of the following statements describe something similar to what Romans meant by res publica?
Choices: [(a) a group of friends that do whatever they want, whenever they want, (b) a class project group where everyone follows the same guidelines and benefits equally from working together, (c) a soccer team where bad players have to do what the coach says, but the best players don’t]
Answer index: 1
Subject: Social Studies
Skill: french-and-english-expeditions-part-2
Description: After John Cabot’s expedition, the English government did not sponsor another expedition to the Americas for more than 50 years. Then, in the mid-1500s, the English government began sending expeditions again. Read the passage about the events in England that created new interest in expeditions. Then answer the question below.
In the middle of the 1500s, England’s main source of wealth was trading English cloth in foreign markets. Thousands of people in England made their livings by producing or selling cloth, and the English government made money by taxing the sale of cloth. But in 1551, some foreign powers stopped allowing people to purchase English cloth. Many English people who worked in the cloth industry lost their jobs, and the government could not raise as much money.
These conditions created a national crisis. The Queen of England, Elizabeth I, needed to find new sources of wealth to address the crisis. Queen Elizabeth thought that sending expeditions to the Americas might be one way to help.
Based on the passage, why did Queen Elizabeth I decide to begin sending expeditions to the Americas?
Choices: [(a) She wanted to stop Spain from purchasing English cloth., (b) She needed to find a new way to make money for England after the cloth crisis., (c) She was curious to learn more about what the Americas looked like.]
Answer index: 1
Subject: Social Studies
Skill: comparing-athens-and-sparta-part
@@
Description: Although city-states such as Athens could trade for food, Sparta was surrounded by mountains that made trade difficult. So, the Spartans relied on agriculture. The ancient Greek historian Pausanias described the role of helots, or enslaved people, in Spartan agriculture. Read the passage. Then complete the text below.
The helots had to bring half of all the crops they grew to the Spartans. Helots who disobeyed were severely punished. The poet Tyrtaeus compared the helots to donkeys exhausted by a heavy load, under painful duty to bring their Spartan masters half of the crops from the land they ploughed and cared for.
Adapted from Pausanias, Description of Greece, translated by W.H.S. Jones and H.A. Ormerod.
Complete the text.
The helots were enslaved people in Sparta. The Spartans relied on the helots ___.
Choices: [(a) to educate their children, (b) to farm their land and raise crops, (c) to trade what was farmed, (d) to train their children to farm]
Answer index: 1
Subject: Social Studies
Skill: thomas-edison
Description: Several inventors worked on lightbulbs before Thomas Edison did. How were Thomas Edison’s lightbulbs different?
@@
Choices: [(a) His lightbulbs were free., (b) His lightbulbs were made from all recycled materials., (c) His lightbulbs could last for several weeks or months., (d) His lightbulbs came in different colors.]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-turning-the-tide-of-war
Description: Both the Continental Army and people on the homefront had trouble paying for the things that they needed during the war. Why?
@@
Choices: [(a) Inflation caused prices to rise., (b) The Continental Congress stopped printing Continentals., (c) The British took gold reserves from the Thirteen Colonies., (d) Most of the money was spent on military supplies.]
Answer index: 0
@@
Subject: Social Studies
Skill: foundations-of-maya-civilization
Description: Complete the text.
@@
Archaeologists believe that the ancestors of the Maya first arrived in Central America more than 12,000 years ago. At first, these people were ___ and traveled long distances to find food. But over time, they learned to grow crops instead. By at least 1500 BCE, they began to settle in permanent ___. This was the beginning of the Maya civilization.
Choices: [(a) explorers . . . cities, (b) farmers . . . palaces, (c) hunter-gatherers . . . villages]
Answer index: 2
Subject: Social Studies
Skill: classical-athens-geography-and-society
Description: Complete the text.
For most of its history, ancient Athens ruled itself and was not part of a larger country. In other words, Athens was a ___.
Choices: [(a) region, (b) city-state, (c) kingdom]
Answer index: 1
Subject: Social Studies
Skill: the-mongol-empire
Description: The founder of the Mongol Empire was a man named Temujin. He was born in Mongolia in the 1100s. The passage below comes from a Mongolian history text. In the passage, an older person is telling Temujin what Mongolia was like before he was born. Read the passage. Then answer the question below.
Before you were born, the stars turned in the heavens. Everyone was fighting. Instead of sleeping, they robbed each other of their possessions. . . . Instead of resting, they fought each other. In such a world one did not live as one wished, but rather in constant conflict. There was no rest, only battle. There was no affection, only mutual slaughter.
Adapted from The Secret History of the Mongols, paragraph 254
Based on the passage, which choice best summarizes what Mongolian society was like in the 1100s?
Choices: [(a) Mongolian society was peaceful, but foreign invaders often attacked Mongolia. As a result, Mongols were often forced to fight, and many Mongols died., (b) People in Mongolia often fought wars with China, their neighbor. As a result, people often went to China to fight., (c) People in Mongolia were fighting with one another and stealing each others’ goods. As a result, people could not live peaceful lives.]
Answer index: 2
Subject: Social Studies
Skill: foundations-of-aztec-civilization
Description: Today, the part of Central America where the Aztec Empire existed is known as Mesoamerica. There are two ways of defining a region like Mesoamerica:
using a geographic definition, which focuses on an area of land and its features, such as mountain ranges and bodies of water
using a cultural definition, which focuses on the customs and practices that people living there share
Which of the following statements uses a geographic definition of Mesoamerica?
Choices: [(a) Mesoamerica is a place where people used a 260-day calendar., (b) Mesoamerica is a place in Central America where people built religious temples in the form of pyramids., (c) Mesoamerica is a region of Central America that includes the Valley of Mexico and the Yucatan Peninsula.]
Answer index: 2
Subject: Language Arts
Skill: choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
Description: Pick the uppercase letter that matches.
k
Choices: [(a) T, (b) R, (c) K]
Answer index: 2
Subject: Language Arts
Skill: choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
Description: Pick the lowercase letter that matches.
C
Choices: [(a) c, (b) v, (c) o]
Answer index: 0
Subject: Language Arts
Skill: choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r
Description: Pick the uppercase letter that matches. a
Choices: [(a) N, (b) A]
Answer index: 1
Subject: Language Arts
Skill: choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r
Description: Pick the lowercase letter that matches. A
Choices: [(a) u, (b) a, (c) n]
Answer index: 1
Subject: Language Arts
Skill: choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y
Description: Pick the lowercase letter that matches.
L
Choices: [(a) l, (b) k, (c) m]
Answer index: 0
Subject: Language Arts
Skill: choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y
Description: Pick the uppercase letter that matches.
m
Choices: [(a) H, (b) J, (c) M]
Answer index: 2
Subject: Language Arts
Skill: distinguish-characters-points-of-view
Description: The following texts both describe a school bus ride on the first day of school.
This text is told from the point of view of Jeffrey, a student going to a new school:
@@
I felt ready for my first shift as a school bus driver. I turned the long bus at each corner with ease. My cheerful mood made the students smile. I was proud of myself as the bus approached the brick school building. My first trip had gone well!
How is Ms. Romero’s point of view different from Jeffrey’s?
Choices: [(a) Ms. Romero describes feeling uneasy about a new experience., (b) Ms. Romero expresses a feeling of pride about the first day of school.]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-words-with-greek-and-latin-roots
Description: What does the word autopilot mean?
@@
Choices: [(a) a tool that pilots use when they want to make the airplane fly faster, (b) a device that flies an airplane on its own, without a pilot, (c) a map or drawing that shows an airplane’s route]
Answer index: 1
Subject: Language Arts
Skill: describe-the-difference-between-related-words
Description: What is the difference between being vicious and being mean?
Choices: [(a) being vicious is worse, (b) being mean shows more anger]
Answer index: 0
Subject: Language Arts
Skill: which-word-does-not-rhyme
Description: Which word does not rhyme?
Choices: [(a) pack, (b) pick, (c) sick]
Answer index: 0
Subject: Language Arts
Skill: complete-the-sentence-with-the-correct-sight-word
Description: Complete the sentence.
___ pig is big.
Choices: [(a) That, (b) Went]
Answer index: 0
Subject: Language Arts
Skill: to-have-use-the-correct-form
Description: Complete the sentence.
Mrs. Barton’s rug ___ blue flowers on it.
Choices: [(a) has, (b) have]
Answer index: 0
Subject: Language Arts
Skill: complete-the-sentence-with-a-two-syllable-word
Description: Which word makes more sense in the sentence?
We ride in the ___
Choices: [(a) wagon, (b) robin]
Answer index: 0
Subject: Language Arts
Skill: complete-the-sentence-with-a-three-letter-consonant-blend-word
Description: Which word makes more sense in the sentence?
Look, that big fish is swimming up the ___
Choices: [(a) stream, (b) strict]
Answer index: 0
Subject: Language Arts
Skill: use-subordinating-conjunctions
Description: Complete the sentence with the better conjunction.
___ Joseph arrives soon, we won’t get to the movie on time.
Choices: [(a) Unless, (b) After]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-homophone
Description: Complete the sentence with the correct homophone.
@@
Given that there were more than one hundred guests at the wedding, how many ___ do you think Martina and Cooper have received?
Choices: [(a) presence, (b) presents]
Answer index: 1
Subject: Language Arts
Skill: use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
Description: The word millimeter contains the root mille. What does the root mille mean?
Choices: [(a) measure, (b) one hundred, (c) one thousand]
Answer index: 2
Subject: Language Arts
Skill: analogies
Description: Complete the analogy.
Ohio is to United States as Earth is to
Choices: [(a) moon, (b) solar system]
Answer index: 1
@@
Subject: Language Arts
Skill: is-it-a-complete-sentence-or-a-fragment
Description: Is this a complete sentence?
Monica speaks four languages.
Choices: [(a) Monica speaks four languages.]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-phrase-or-a-clause
Description: Is the group of words in bold a phrase or a clause?
*Emily’s suitcase was too large to fit in the overhead bin**, so unfortunately she had to check it.
Choices: [(a) phrase, (b) clause]
Answer index: 1
Subject: Language Arts
Skill: choose-between-subject-and-object-pronouns
Description: Select the correct pronoun to complete the sentence.
@@
"Thanks to your efforts, ___ have emerged as leaders in the energy industry," Mr. Bullock announced at the company meeting.
Choices: [(a) us, (b) we]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-simple-compound-complex-or-compound-complex
Description: Which type of sentence is this?